OpenStudy (anonymous):

Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7 @KailaMorgan7777

OpenStudy (anonymous):

@KailaMorgan7777

OpenStudy (anonymous):

question 3 out of 5 and m latop is about to die coukd you stay adn answer 2 more?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

What is the solution to the equation? x – 1.2 = 3.6 x =

OpenStudy (anonymous):

opps wait ummm let me just take a scrren shot of that

OpenStudy (anonymous):

3.6+1.2=?

OpenStudy (anonymous):

x – 1.2 = 3.6 what is x?

OpenStudy (anonymous):

then whatever you get is what x is

OpenStudy (anonymous):

kk alt one that i need to ask is this one let me get it

OpenStudy (anonymous):

Use a transformation to solve the equation. s – 14.2 = –18.7 A. s = –32.9 B. s = –4.5 C. s = 4.5 D. s = 32.9

OpenStudy (anonymous):